The difference of two numbers is eleven. The second is three less than two times the first. what are the two numbers?
Answer by checkley71(8403)

X-Y=11
Y=2X-3
THEN X=11+Y
Y=2(11+Y)-3
Y=22+2Y-3
Y-2Y=22-3
-Y=19
Y=-19
-19=2X-3
2X=-19+3
2X=-16
X=-8
PROOF
-8-(-19=11
-8+19=11
11=11